# No-go theorem for probabilistic one-way secret-key distillation

### v1

@@ Appendix A Equivalent definitions of probabilistic one-way distillable secret key
where $\gamma^{k}_{A^{\prime}B^{\prime}}$ is a bipartite private state holding $\log_{2}k$ secret key bits and $\sigma_{A^{\prime}B^{\prime}}$ is an arbitrary bipartite state. As such, this protocol can be used to distill an expected number $p\log_{2}k$ of secret key bits from a single instance of $\rho_{AB}$. The one-shot probabilitic one-way distillable secret key of the state $\rho_{AB}$ is defined generally as
$K_{D}^{(1),\to}(\rho_{AB})\coloneqq\\ \sup_{\begin{subarray}{c}p\in[0,1],k\in\mathbb{N}\\ \mathcal{L}^{\to}\in\operatorname{1WL},\gamma^{k}_{A^{\prime}B^{\prime}}\end{subarray}}\left\{\begin{array}[]{c}p\log_{2}k:\\ \mathcal{L}^{\to}\!\left(\rho_{AB}\right)=p[1]_{X_{A}}\otimes\gamma^{k}_{A^{\prime}B^{\prime}}\\ +(1-p)[0]_{X_{A}}\otimes\sigma_{A^{\prime}B^{\prime}},\\ \sigma_{A^{\prime}B^{\prime}}\in\mathcal{S}\!\left(A^{\prime}B^{\prime}\right)\end{array}\right\}.$ (16)
@@
$(\mathcal{C}_{X_{A}\to X_{A}X_{B}}\circ\mathcal{L}^{\to})\left(\rho_{AB}\right)=p[1]_{X_{A}}\otimes[1]_{X_{B}}\otimes\gamma^{k}_{A^{\prime}B^{\prime}}\\ +(1-p)[0]_{X_{A}}\otimes[0]_{X_{B}}\otimes\sigma_{A^{\prime}B^{\prime}},$ (17)
@@
$\mathcal{R}_{X_{A}A^{\prime}\to X_{A}A^{\prime}}\!\left(\cdot\right)=\frac{I_{A^{\prime}}}{d_{A^{\prime}}}\otimes\operatorname{Tr}_{B^{\prime}}\!\left[[0]_{X_{B}}\left(\cdot\right)[0]_{X_{B}}\right]\\ +\operatorname{id}_{X_{A}A^{\prime}\to X_{A}A^{\prime}}\!\left([1]_{X_{A}}\left(\cdot\right)[1]_{X_{A}}\right),$ (18)
@@
$(\mathcal{R}\circ\mathcal{C}\circ\mathcal{L}^{\to})\!\left(\rho_{AB}\right)$
$=p[1]_{X_{A}}\otimes\gamma^{k}_{A^{\prime}B^{\prime}}+(1-p)[0]_{X_{A}}\otimes\frac{I_{A^{\prime}}}{d_{A^{\prime}}}\otimes\frac{I_{B^{\prime}}}{d_{B^{\prime}}}$ (19)
@@
Observe that $\mathcal{R}\circ\mathcal{C}\circ\mathcal{L}^{\to}$ is also a one-way LOCC channel that distills an expected number $p\log_{2}k$ of secret key bits from a single instance of $\rho_{AB}$. Since such a one-way LOCC transformation can be designed for every one-way LOCC channel $\mathcal{L}^{\to}_{AB\to X_{A}A^{\prime}B^{\prime}}$, we can conclude the following statement: if there exists a one-way LOCC channel that distills a quantum state of the form given in (15), then there also exists a one-way LOCC channel that distills the quantum state of the form given in (20). As the expected number of secret key bits distilled from $\rho_{AB}$ using either of the channels is equal to $p\log_{2}k$, we can restrict the optimization in the definition of one-shot probabilistic one-way distillable secret key to channels that establish a quantum state of the form given in (20). Hence, we arrive at the following equivalent definition of the one-shot probabilistic one-way distillable secret key given in (5).
@@ Appendix B The min-unextendible entanglement of bipartite states
The unextendible entanglement provides a framework to quantify the unextendibility of a bipartite state $\rho_{AB}$ with respect to the system $B$. Crucially, this quantity is monotonic under the action of two-extendible channels [22, Theorem 2], and hence, it is also monotonic under the action of one-way LOCC channels. (See [20, 21, 22] for the definition of a two-extendible channel.) Stated formally, [22, Theorem 2] establishes the following: for every quantum state $\rho_{AB}$ and two-extendible channel $\mathcal{N}_{AB\to A^{\prime}B^{\prime}}$, the following inequality holds:
@@
A different measure for unextendibility was considered in [20, 21] where the divergence was measured with respect to a fixed set of two-extendible states. However, in Definition 2, the divergence is measured by means of a set of states that depend on the input state itself. Although both measures are equal to the minimal possible value of $\mathbf{D}$ when $\rho_{AB}$ is two-extendible, they are not equal in general.
@@
where $D_{\min}$ is the min-relative entropy [23, Def. 2]:
@@
    Non-negativity: The min-unextendible entanglement is non-negative for a bipartite quantum state, as a consequence of the min-relative entropy being non-negative for all pairs of states; that is,
    $E^{u}_{\min}\!\left(\rho_{AB}\right)\geq 0\quad\forall\rho_{AB}\in\mathcal{S}\!\left(AB\right).$ (25)
@@
    Additivity [22, Prop. 15]: The min-unextendible entanglement is additive with respect to a tensor product of the states $\rho_{A_{1}B_{1}}$ and $\sigma_{A_{2}B_{2}}$:
@@
    Privacy bound [22, Prop. 21] The min-unextendible entanglement of a bipartite private state $\gamma^{k}_{AB}$, holding $\log_{2}k$ secret key bits, is bounded from below as follows:
@@ Appendix C Proof of Lemma 1
Consider the following dephasing channel:
@@
$\eta^{p,k}_{AB}=p~{}\gamma^{k}_{AB}+(1-p)[e]_{A}\otimes[e]_{B},$ (34)
with $\gamma^{k}_{AB}$ a bipartite private state holding $\log_{2}k$ secret key bits. Here systems $E$ and $B$ are isomorphic to each other. The action of the dephasing channel $\Delta_{A}$, defined in (28), on the state $\omega_{AE}$ results in the following state:
$\Delta_{A}\!\left(\omega_{AE}\right)=p~{}\sigma_{AE}+(1-p)[e]_{A}\otimes\tau_{E},$ (35)
@@
Recall the definition of the dephasing channel $\Delta_{A}$ from (28). The state arising from the action of $\Delta_{A}$ on the state $\omega_{ABE}$ can be decomposed as follows:
@@
where $[e]_{A}X_{ABE}=X_{ABE}[e]_{A}=0$. Since $\Delta_{A}\!\left(\omega_{ABE}\right)$ is a quantum state, $\Delta\!\left(\omega_{ABE}\right)\geq 0$, which implies the following operator inequalities:
@@
where we have used the fact that $X_{ABE}$ and $[e]_{A}\otimes Y_{BE}$ are orthogonal to each other.
@@
$\operatorname{Tr}_{E}\!\left[\Delta_{A}\!\left(\omega_{ABE}\right)\right]$ $=\operatorname{Tr}_{E}\!\left[X_{ABE}\right]+\operatorname{Tr}_{E}\!\left[Y_{BE}\right]\otimes[e]_{A}$ (42)
We can also evaluate the marginal by first tracing out the system $E$ and then applying the dephasing channel; that is,
@@
$=p~{}\gamma^{k}_{AB}+(1-p)[e]_{A}\otimes[e]_{B}.$ (44)
Comparing (42) and (44), and using the fact that $[e]_{A}X_{ABE}=0$, we arrive at the following equalities:
$\operatorname{Tr}_{E}\!\left[X_{ABE}\right]$ $=p~{}\gamma^{k}_{AB},$ (45)
@@
Recall from (39) and (41) that $X_{ABE}$ and $Y_{BE}$ are positive semidefinite operators. Using (45), we can further conclude that $\frac{X_{ABE}}{p}$ is a quantum state that extends $\gamma^{k}_{AB}$. Similarly, (46) implies that $\frac{Y_{BE}}{1-p}$ is a quantum state that extends $[e]_{B}$.
Any quantum state that extends $[e]_{B}$ is of the form $[e]_{B}\otimes\tau_{E}$, where $\tau_{E}$ is an arbitrary quantum state (and which can even be $[e]_{E}$). Let us define the quantum state $\sigma_{AE}\coloneqq\frac{1}{p}\operatorname{Tr}_{B}\!\left[X_{ABE}\right]$. Since $\frac{1}{p}\operatorname{Tr}_{E}\!\left[X_{ABE}\right]=\gamma^{k}_{AB}$, $\sigma_{AE}$ is in the set $\mathcal{F}\!\left(\gamma^{k}_{AB}\right)$. The marginal of the state $\Delta_{A}\!\left(\omega_{ABE}\right)$ on systems $AE$ can now be written as follows:
$\Delta_{A}\!\left(\omega_{AE}\right)=p~{}\sigma_{AE}+(1-p)[e]_{A}\otimes\tau_{E},$ (47)
@@
Proof of Lemma 1. The min-unextendible entanglement of $\eta^{p,k}_{AB}$ is as follows:
@@
where $\eta^{p,k}_{AB}=p~{}\gamma^{k}_{AB}+(1-p)[e]_{A}\otimes[e]_{B}$ with $\gamma^{k}_{AB}$ a bipartite private state holding $\log_{2}k$ secret key bits.
@@
where $\Delta_{A}$ is defined in (28). The above inequality follows from the data-processing of min-relative entropy and the subsequent equality follows from the invariance of the doubly erased state under the action of the dephasing channel $\Delta_{A}$. We have shown in Lemma 2 that $\Delta_{A}\!\left(\omega_{AB}\right)$ also is in the set $\mathcal{F}\!\left(\eta^{p,k}_{AB}\right)$. Then the inequality in (49)–(50) implies that we can restrict the optimization in (48) to states $\omega_{AB}$ such that $\omega_{AB}=\Delta_{A}\!\left(\omega_{AB}\right)$. Alternatively, we can write the min-unextendible entanglement of the doubly erased state as follows:
@@
$\Delta_{A}\!\left(\omega_{AB}\right)=p~{}\sigma_{AB}+(1-p)[e]_{A}\otimes\tau_{B},$ (52)
@@
$E^{u}_{\min}\!\left(\eta^{p,k}_{AB}\right)$
$=\inf_{\omega_{AB}\in\mathcal{F}\left(\eta^{p,k}\right)}\frac{1}{2}D_{\min}\!\left(\eta^{p,k}_{AB}\Big{\|}\Delta_{A}\!\left(\omega_{AB}\right)\right)$ (53)
$=\inf_{\sigma,\tau}-\frac{1}{2}\log_{2}\operatorname{Tr}\!\left[\Pi^{\eta^{p,k}}\left(p~{}\sigma_{AB}+(1-p)[e]_{A}\otimes\tau_{B}\right)\right]$ (54)
$=-\frac{1}{2}\log_{2}\sup_{\sigma,\tau}\operatorname{Tr}\!\left[\Pi^{\eta^{p,k}}\left(p~{}\sigma_{AB}+(1-p)[e]_{A}\otimes\tau_{B}\right)\right],$ (55)
where $\Pi^{\eta^{p,k}}$ is the projection onto the support of $\eta^{p,k}_{AB}$. The optimization in the second and third equality is over all $\sigma_{AB}\in\mathcal{F}\!\left(\gamma^{k}_{AB}\right)$ and $\tau_{B}\in\mathcal{S}\!\left(B\right)$. The last equality follows from the monotonicity of the logarithm.
Note that $\gamma^{k}_{AB}$ is orthogonal to $[e]_{A}\otimes[e]_{B}$, and $[e]_{A}\otimes[e]_{B}$ is a projection. Therefore, we can write the projection onto the support of $\eta^{p,k}_{AB}$ as the following sum:
@@
$E^{u}_{\min}\!\left(\eta^{p,k}_{AB}\right)$
$=-\frac{1}{2}\log_{2}\sup_{\sigma,\tau}\Big{(}p\operatorname{Tr}\!\left[\Pi^{\gamma^{k}}_{AB}\sigma_{AB}\right]$
$\qquad\qquad+(1-p)\operatorname{Tr}\!\left[\left([e]_{A}\otimes[e]_{B}\right)\left([e]_{A}\otimes\tau_{B}\right)\right]\Big{)}$ (57)
$=-\frac{1}{2}\log_{2}\sup_{\sigma,\tau}\Big{(}p\operatorname{Tr}\!\left[\Pi^{\gamma^{k}}_{AB}\sigma_{AB}\right]$
$\qquad\qquad\qquad\qquad+(1-p)\operatorname{Tr}\!\left[[e]_{A}\otimes\left([e]_{E}\tau_{B}\right)\right]\Big{)},$ (58)
where the first equality follows from the fact that $\left([e]_{A}\otimes I_{B}\right)\sigma_{AB}=0$ and $\Pi^{\gamma^{k}}_{AB}\left([e]_{A}\otimes\tau_{B}\right)=0$ for every quantum state $\tau_{E}$.
It is clear that $\tau_{B}$ should be $[e]_{B}$ in order for the second term inside the logarithm in (58) to be non-zero. Therefore,
$E^{u}_{\min}\!\left(\eta^{p,k}_{AB}\right)$
$=-\frac{1}{2}\log_{2}\sup_{\sigma\in\mathcal{F}\left(\gamma^{k}_{AB}\right)}\left(p\operatorname{Tr}\!\left[\Pi^{\gamma^{k}}\sigma_{AB}\right]+1-p\right)$ (59)
$=-\frac{1}{2}\log_{2}\sup_{\sigma\in\mathcal{F}\left(\gamma^{k}_{AB}\right)}\left(p~{}2^{-D_{\min}\!\left(\gamma^{k}_{AB}\Big{\|}\sigma_{AB}\right)}+1-p\right)$ (60)
$=-\frac{1}{2}\log_{2}\left(p~{}2^{-\inf_{\sigma\in\mathcal{F}\left(\gamma^{k}_{AB}\right)}D_{\min}\!\left(\gamma^{k}_{AB}\Big{\|}\sigma_{AB}\right)}+1-p\right)$ (61)
$=-\frac{1}{2}\log_{2}\left(p~{}2^{-2E^{u}_{\min}\!\left(\gamma^{k}_{AB}\right)}+1-p\right),$ (62)
@@
It has been shown in [22, Corollary 22] that the min-unextendible entanglement of a bipartite private state holding $\log_{2}k$ bits of secrecy is not less than $\log_{2}k$; that is,
@@
$p~{}2^{-2E^{u}_{\min}\!\left(\gamma^{k}\right)}\leq p~{}2^{-2\log_{2}k}=\frac{p}{k^{2}}.$ (64)
@@
$\log_{2}\left(p~{}2^{-\inf_{\sigma}D_{\min}\!\left(\gamma^{k}_{AB}\Big{\|}\operatorname{Tr}_{B}\!\left[\sigma_{ABE}\right]\right)}+1-p\right)\\ \leq\log_{2}\left(\frac{p}{k^{2}}+1-p\right),$ (65)
@@
$E^{u}_{\min}\!\left(\eta^{p,k}_{AB}\right)\geq-\frac{1}{2}\log_{2}\left(\frac{p}{k^{2}}+1-p\right)\quad\forall p\in(0,1),k\geq 2.$ (66)
When $p=0$, $\eta^{p,k}_{AB}$ is a separable state, and hence, its min-unextendible entanglement is equal to zero [22, Proposition 3]. When $p=1$, $\eta^{p,k}_{AB}$ is a bipartite private state holding $\log_{2}k$ bits of secrecy, and hence, its min-uenxtendible entanglement is not less than $\log_{2}k$ as stated in [22, Corollary 22]. Thus, we conclude the statement of Lemma 1.
Appendix D Convexity and openness of the set of super two-extendible states

### v2

@@ Appendix A Equivalent definitions of probabilistic one-way distillable secret key
where $\gamma^{k}_{A^{\prime}B^{\prime}}$ is a bipartite private state holding $\log_{2}k$ secret key bits and $\sigma_{A^{\prime}B^{\prime}}$ is an arbitrary bipartite state. As such, this protocol can be used to distill an expected number $p\log_{2}k$ of secret key bits from a single instance of $\rho_{AB}$. The one-shot probabilistic one-way distillable secret key of the state $\rho_{AB}$ is defined generally as
$K_{D}^{(1),\to}(\rho_{AB})\coloneqq\sup_{\begin{subarray}{c}p\in[0,1],k\in\mathbb{N}\\ \mathcal{L}^{\to}\in\operatorname{1WL},\gamma^{k}_{A^{\prime}B^{\prime}}\end{subarray}}\left\{\begin{array}[]{c}p\log_{2}k:\\ \mathcal{L}^{\to}\!\left(\rho_{AB}\right)=p[1]_{X_{A}}\otimes\gamma^{k}_{A^{\prime}B^{\prime}}\\ +(1-p)[0]_{X_{A}}\otimes\sigma_{A^{\prime}B^{\prime}},\\ \sigma_{A^{\prime}B^{\prime}}\in\mathcal{S}\!\left(A^{\prime}B^{\prime}\right)\end{array}\right\}.$ (18)
@@
$(\mathcal{C}_{X_{A}\to X_{A}X_{B}}\circ\mathcal{L}^{\to})\left(\rho_{AB}\right)=p[1]_{X_{A}}\otimes[1]_{X_{B}}\otimes\gamma^{k}_{A^{\prime}B^{\prime}}+(1-p)[0]_{X_{A}}\otimes[0]_{X_{B}}\otimes\sigma_{A^{\prime}B^{\prime}},$ (19)
@@
$\mathcal{R}_{X_{A}A^{\prime}\to X_{A}A^{\prime}}\!\left(\cdot\right)=\frac{I_{A^{\prime}}}{d_{A^{\prime}}}\otimes\operatorname{Tr}_{A^{\prime}}\!\left[[0]_{X_{A}}\left(\cdot\right)[0]_{X_{A}}\right]+\operatorname{id}_{X_{A}A^{\prime}\to X_{A}A^{\prime}}\!\left([1]_{X_{A}}\left(\cdot\right)[1]_{X_{A}}\right),$ (20)
@@
$(\mathcal{R}\circ\mathcal{C}\circ\mathcal{L}^{\to})\!\left(\rho_{AB}\right)$ $=p[1]_{X_{A}}\otimes\gamma^{k}_{A^{\prime}B^{\prime}}+(1-p)[0]_{X_{A}}\otimes\frac{I_{A^{\prime}}}{d_{A^{\prime}}}\otimes\frac{I_{B^{\prime}}}{d_{B^{\prime}}}$ (21)
@@
Observe that $\mathcal{R}\circ\mathcal{C}\circ\mathcal{L}^{\to}$ is also a one-way LOCC channel that distills an expected number $p\log_{2}k$ of secret key bits from a single instance of $\rho_{AB}$. Since such a one-way LOCC transformation can be designed for every one-way LOCC channel $\mathcal{L}^{\to}_{AB\to X_{A}A^{\prime}B^{\prime}}$, we can conclude the following statement: if there exists a one-way LOCC channel that distills a quantum state of the form given in (17), then there also exists a one-way LOCC channel that distills the quantum state of the form given in (22). As the expected number of secret key bits distilled from $\rho_{AB}$ using either of the channels is equal to $p\log_{2}k$, we can restrict the optimization in the definition of one-shot probabilistic one-way distillable secret key to channels that establish a quantum state of the form given in (22). Hence, we arrive at the equivalent definition of the one-shot probabilistic one-way distillable secret key given in (6).
@@ Appendix B The min-unextendible entanglement of bipartite states
The unextendible entanglement provides a framework for quantifying the unextendibility of a bipartite state $\rho_{AB}$ with respect to the system $B$. Crucially, this quantity is monotonic under the action of two-extendible channels [25, Theorem 2], and hence, it is also monotonic under the action of one-way LOCC channels. (See [23, 24, 25] for the definition of a two-extendible channel.) Stated formally, [25, Theorem 2] establishes the following: for every quantum state $\rho_{AB}$ and two-extendible channel $\mathcal{N}_{AB\to A^{\prime}B^{\prime}}$, the following inequality holds:
@@
A different measure for unextendibility was proposed in [23, 24] where the divergence was measured with respect to a fixed set of two-extendible states. However, in Definition 2, the divergence is measured by means of a set of states that depend on the input state itself. Although both measures are equal to the minimal possible value of $\mathbf{D}$ when $\rho_{AB}$ is two-extendible, they are not equal in general.
@@
where $D_{\min}$ is the min-relative entropy [26, Def. 2]:
@@
    Non-negativity: The min-unextendible entanglement is non-negative for a bipartite quantum state, as a consequence of the min-relative entropy being non-negative for all pairs of states. That is,
    $E^{u}_{\min}\!\left(\rho_{AB}\right)\geq 0\qquad\forall\rho_{AB}\in\mathcal{S}\!\left(AB\right).$ (27)
@@
    Additivity [25, Prop. 15]: The min-unextendible entanglement is additive with respect to a tensor product of the states $\rho_{A_{1}B_{1}}$ and $\sigma_{A_{2}B_{2}}$:
@@
    Privacy bound [25, Prop. 21] The min-unextendible entanglement of a bipartite private state $\gamma^{k}_{AB}$, holding $\log_{2}k$ secret key bits, is bounded from below as follows:
@@ Appendix C Proof of Lemma 1
We restate Lemma 1 here for the reader’s convenience.
Consider the following partially dephasing channel:
@@
$\eta^{p,k}_{AB}=p\penalty 10000\ \gamma^{k}_{AB}+(1-p)[e]_{A}\otimes[e]_{B},$ (37)
with $\gamma^{k}_{AB}$ a bipartite private state holding $\log_{2}k$ secret key bits. Here systems $E$ and $B$ are isomorphic to each other. The action of the partially dephasing channel $\Delta_{A}$, defined in (31), on the state $\omega_{AE}$ results in the following state:
$\Delta_{A}\!\left(\omega_{AE}\right)=p\penalty 10000\ \sigma_{AE}+(1-p)[e]_{A}\otimes\tau_{E},$ (38)
@@
Recall the definition of the partially dephasing channel $\Delta_{A}$ from (31). The state arising from the action of $\Delta_{A}$ on the state $\omega_{ABE}$ can be decomposed as follows:
@@
where $e]_{A}X_{ABE}=X_{ABE}[e]_{A}=0$. Since $\Delta_{A}\!\left(\omega_{ABE}\right)$ is a quantum state, $\Delta\!\left(\omega_{ABE}\right)\geq 0$, which implies the following operator inequalities:
@@
where we have used the fact that $X_{ABE}$ and $e]_{A}\otimes Y_{BE}$ are orthogonal to each other.
@@
$\operatorname{Tr}_{E}\!\left[\Delta_{A}\!\left(\omega_{ABE}\right)\right]$ $=\operatorname{Tr}_{E}\!\left[X_{ABE}\right]+\operatorname{Tr}_{E}\!\left[Y_{BE}\right]\otimes[e]_{A}.$ (45)
We can also evaluate the marginal by first tracing out the system $E$ and then applying the partially dephasing channel; that is,
@@
$=p\penalty 10000\ \gamma^{k}_{AB}+(1-p)[e]_{A}\otimes[e]_{B}.$ (47)
Comparing (45) and (47), and using the fact that $e]_{A}X_{ABE}=0$, we arrive at the following equalities:
$\operatorname{Tr}_{E}\!\left[X_{ABE}\right]$ $=p\penalty 10000\ \gamma^{k}_{AB},$ (48)
@@
Recall from (42) and (44) that $X_{ABE}$ and $Y_{BE}$ are positive semidefinite operators. Using (48), we can further conclude that $\frac{X_{ABE}}{p}$ is a quantum state that extends $\gamma^{k}_{AB}$. Similarly, (49) implies that $\frac{Y_{BE}}{1-p}$ is a quantum state that extends $e]_{B}$.
Any quantum state that extends $e]_{B}$ is of the form $e]_{B}\otimes\tau_{E}$, where $\tau_{E}$ is an arbitrary quantum state (and which can even be $e]_{E}$). Let us define the quantum state $\sigma_{AE}\coloneqq\frac{1}{p}\operatorname{Tr}_{B}\!\left[X_{ABE}\right]$. Since $\frac{1}{p}\operatorname{Tr}_{E}\!\left[X_{ABE}\right]=\gamma^{k}_{AB}$, $\sigma_{AE}$ is in the set $\mathcal{F}\!\left(\gamma^{k}_{AB}\right)$. The marginal of the state $\Delta_{A}\!\left(\omega_{ABE}\right)$ on systems $AE$ can now be written as follows:
$\Delta_{A}\!\left(\omega_{AE}\right)=p\penalty 10000\ \sigma_{AE}+(1-p)[e]_{A}\otimes\tau_{E},$ (50)
@@
Proof of Lemma 1. The min-unextendible entanglement of $\eta^{p,k}_{AB}$ is equal to the following:
@@
where $\eta^{p,k}_{AB}=p\penalty 10000\ \gamma^{k}_{AB}+(1-p)[e]_{A}\otimes[e]_{B}$ with $\gamma^{k}_{AB}$ a bipartite private state holding $\log_{2}k$ secret key bits.
@@
where $\Delta_{A}$ is defined in (31). The above inequality follows from the data-processing of min-relative entropy and the subsequent equality follows from the invariance of the doubly erased state under the action of the partially dephasing channel $\Delta_{A}$. We have shown in Lemma 2 that $\Delta_{A}\!\left(\omega_{AB}\right)$ also is in the set $\mathcal{F}\!\left(\eta^{p,k}_{AB}\right)$. Then the inequality in (52)–(53) implies that we can restrict the optimization in (51) to states $\omega_{AB}$ such that $\omega_{AB}=\Delta_{A}\!\left(\omega_{AB}\right)$. Alternatively, we can write the min-unextendible entanglement of the doubly erased state as follows:
@@
$\Delta_{A}\!\left(\omega_{AB}\right)=p\penalty 10000\ \sigma_{AB}+(1-p)[e]_{A}\otimes\tau_{B},$ (55)
@@
$E^{u}_{\min}\!\left(\eta^{p,k}_{AB}\right)$ $=\inf_{\omega_{AB}\in\mathcal{F}\left(\eta^{p,k}\right)}\frac{1}{2}D_{\min}\!\left(\eta^{p,k}_{AB}\Big\|\Delta_{A}\!\left(\omega_{AB}\right)\right)$ (56)
$=\inf_{\sigma,\tau}-\frac{1}{2}\log_{2}\operatorname{Tr}\!\left[\Pi^{\eta^{p,k}}\left(p\penalty 10000\ \sigma_{AB}+(1-p)[e]_{A}\otimes\tau_{B}\right)\right]$ (57)
$=-\frac{1}{2}\log_{2}\sup_{\sigma,\tau}\operatorname{Tr}\!\left[\Pi^{\eta^{p,k}}\left(p\penalty 10000\ \sigma_{AB}+(1-p)[e]_{A}\otimes\tau_{B}\right)\right],$ (58)
where $\Pi^{\eta^{p,k}}$ is the projection onto the support of $\eta^{p,k}_{AB}$. The optimization in the second and third equalities is over all $\sigma_{AB}\in\mathcal{F}\!\left(\gamma^{k}_{AB}\right)$ and $\tau_{B}\in\mathcal{S}\!\left(B\right)$. The last equality follows from the monotonicity of the logarithm.
Note that $\gamma^{k}_{AB}$ is orthogonal to $e]_{A}\otimes[e]_{B}$, and $e]_{A}\otimes[e]_{B}$ is a projection. Therefore, we can write the projection onto the support of $\eta^{p,k}_{AB}$ as the following sum:
@@
$E^{u}_{\min}\!\left(\eta^{p,k}_{AB}\right)$ $=-\frac{1}{2}\log_{2}\sup_{\sigma,\tau}\left(p\operatorname{Tr}\!\left[\Pi^{\gamma^{k}}_{AB}\sigma_{AB}\right]+(1-p)\operatorname{Tr}\!\left[\left([e]_{A}\otimes[e]_{B}\right)\left([e]_{A}\otimes\tau_{B}\right)\right]\right)$ (60)
$=-\frac{1}{2}\log_{2}\sup_{\sigma,\tau}\left(p\operatorname{Tr}\!\left[\Pi^{\gamma^{k}}_{AB}\sigma_{AB}\right]+(1-p)\operatorname{Tr}\!\left[[e]_{A}\otimes\left([e]_{B}\tau_{B}\right)\right]\right),$ (61)
where the first equality follows from the fact that $\left([e]_{A}\otimes I_{B}\right)\sigma_{AB}=0$ and $\Pi^{\gamma^{k}}_{AB}\left([e]_{A}\otimes\tau_{B}\right)=0$ for every quantum state $\tau_{B}$.
It is clear that $\tau_{B}$ should be $e]_{B}$ in order for the second term inside the logarithm in (61) to be non-zero. Therefore,
$E^{u}_{\min}\!\left(\eta^{p,k}_{AB}\right)$ $=-\frac{1}{2}\log_{2}\sup_{\sigma\in\mathcal{F}\left(\gamma^{k}_{AB}\right)}\left(p\operatorname{Tr}\!\left[\Pi^{\gamma^{k}}\sigma_{AB}\right]+1-p\right)$ (62)
$=-\frac{1}{2}\log_{2}\sup_{\sigma\in\mathcal{F}\left(\gamma^{k}_{AB}\right)}\left(p\penalty 10000\ 2^{-D_{\min}\!\left(\gamma^{k}_{AB}\Big\|\sigma_{AB}\right)}+1-p\right)$ (63)
$=-\frac{1}{2}\log_{2}\left(p\penalty 10000\ 2^{-\inf_{\sigma\in\mathcal{F}\left(\gamma^{k}_{AB}\right)}D_{\min}\!\left(\gamma^{k}_{AB}\Big\|\sigma_{AB}\right)}+1-p\right)$ (64)
$=-\frac{1}{2}\log_{2}\left(p\penalty 10000\ 2^{-2E^{u}_{\min}\!\left(\gamma^{k}_{AB}\right)}+1-p\right),$ (65)
@@
It has been shown in [25, Corollary 22] that the min-unextendible entanglement of a bipartite private state holding $\log_{2}k$ bits of secrecy is not less than $\log_{2}k$. That is,
@@
$p\penalty 10000\ 2^{-2E^{u}_{\min}\!\left(\gamma^{k}\right)}\leq p\penalty 10000\ 2^{-2\log_{2}k}=\frac{p}{k^{2}}.$ (67)
@@
$\log_{2}\left(p\penalty 10000\ 2^{-2E^{u}_{\min}\!\left(\gamma^{k}_{AB}\right)}+1-p\right)\leq\log_{2}\left(\frac{p}{k^{2}}+1-p\right),$ (68)
@@
$E^{u}_{\min}\!\left(\eta^{p,k}_{AB}\right)\geq-\frac{1}{2}\log_{2}\left(\frac{p}{k^{2}}+1-p\right)\qquad\forall p\in(0,1),k\geq 2.$ (69)
When $p=0$, $\eta^{p,k}_{AB}$ is a separable state, and hence, its min-unextendible entanglement is equal to zero [25, Proposition 3]. When $p=1$, $\eta^{p,k}_{AB}$ is a bipartite private state holding $\log_{2}k$ bits of secrecy, and hence, its min-unextendible entanglement is not less than $\log_{2}k$ as stated in [25, Corollary 22]. Thus, we conclude the statement of Lemma 1.
Appendix D The set of super two-extendible states is convex but not closed